# Modeling Cell Dynamics and Interactions with Unbalanced Mean Field Schrödinger Bridge

### v2

@@ Abstract
Modeling the dynamics from sparsely time-resolved snapshot data is crucial for understanding complex cellular processes and behavior. Existing methods leverage optimal transport, Schrödinger bridge theory, or their variants to simultaneously infer stochastic, unbalanced dynamics from snapshot data. However, these approaches remain limited in their ability to account for cell-cell interactions. This integration is essential in real-world scenarios since intercellular communications are fundamental life processes and can influence cell state-transition dynamics. To address this challenge, we formulate the Unbalanced Mean-Field Schrödinger Bridge (UMFSB) framework to model unbalanced stochastic interaction dynamics from snapshot data. Inspired by this framework, we further propose CytoBridge, a deep learning algorithm designed to approximate the UMFSB problem. By explicitly modeling cellular transitions, proliferation, and interactions through neural networks, CytoBridge offers the flexibility to learn these processes directly from data. The effectiveness of our method has been extensively validated using both synthetic gene regulatory data and real scRNA-seq datasets. Compared to existing methods, CytoBridge identifies growth, transition, and interaction patterns, eliminates false transitions, and reconstructs the developmental landscape with greater accuracy.
@@ 1 Introduction
Reconstructing dynamics from high-dimensional distribution samples is a central challenge in science and machine learning. In generative models, methods such as Variational Autoencoders (VAEs) (Kingma & Welling, 2013), diffusion models (Ho et al., 2020; Sohl-Dickstein et al., 2015; Song et al., 2021), and flow matching (Lipman et al., 2023; Tong et al., 2024a) have achieved success in generating high-fidelity images by coupling high-dimensional distributions (Liu et al., 2023). Meanwhile, in biology, inferring dynamics (also known as trajectory inference problem) from several static snapshots of single-cell RNA sequencing (scRNA-seq) data (Ding et al., 2022) to build the continuous dynamics of an individual cell and construct the corresponding cell-fate landscapes has also attracted broad interests (Schiebinger et al., 2019; Klein et al., 2025; Zhang et al., 2025b).
To study the trajectory inference problem, optimal transport (OT) theory serves as a foundational tool (Bunne et al., 2024; Zhang et al., 2025b; Heitz et al., 2024). In particular, several works propose to infer continuous cellular dynamics over time by employing the Benamou–Brenier formulation (Benamou & Brenier, 2000). Given that cell growth and death are critical biological processes, modeling the coupling of underlying unnormalized distributions has led to the development of unbalanced dynamical OT by introducing Wasserstein–Fisher–Rao metric (Chizat et al., 2018a, b). To further account for the prevalent stochastic effects on single-cell level, methods inspired by the Schrödinger Bridge (SB) problem seek to identify the most likely stochastic transition path between two arbitrary distributions (Gentil et al., 2017; Léonard, 2014). To tackle both stochastic and unbalanced effects simultaneously, methods have been developed to model unbalanced stochastic dynamics (Pariset et al., 2023; Lavenant et al., 2024), along with a recent deep learning method (Zhang et al., 2025a), which leverages the regularized unbalanced optimal transport (RUOT) framework to infer continuous unbalanced stochastic dynamics from samples without requiring prior knowledge.
@@ 2 Related Works
Numerous efforts have been made to learn dynamics from snapshot data. To tackle the dynamical optimal transport problem, several methods have been proposed by leveraging neural ODEs or flow matching techniques (Tong et al., 2020; Huguet et al., 2022; Wan et al., 2023; Zhang et al., 2024a; Tong et al., 2024a; Albergo et al., 2023; Palma et al., 2025; Rohbeck et al., 2025; Petrović et al., 2025). To account for sink and source terms in unnormalized distributions, (Peng et al., 2024; Sha et al., 2024; Tong et al., 2023; Eyring et al., 2024) developed the neural ODE-based solver for unbalanced dynamical OT. For the Schrödinger Bridge (SB) problem, approaches have been proposed based on its static or dynamic formulations (Shi et al., 2024; De Bortoli et al., 2021; Gu et al., 2025; Koshizuka & Sato, 2023; Neklyudov et al., 2023, 2024; Zhang et al., 2024b; Bunne et al., 2023; Chen et al., 2022a; Zhou et al., 2024a; Zhu et al., 2024; Maddu et al., 2024; Yeo et al., 2021; Jiang & Wan, 2024), with corresponding flow matching methods (Tong et al., 2024b). To further incorporate unbalanced effects in the SB framework, methods utilizing branching SDE theory (Lavenant et al., 2024; Ventre et al., 2023; Chizat et al., 2022), forward-backward SDE (Pariset et al., 2023), and neural ODEs with Fisher information regularization have been introduced (Zhang et al., 2025a). However, theoretical formulation, along with an effective deep learning solver to simultaneously account for unbalanced stochastic effects and particle interactions in the dynamical OT framework, remains largely lacking.
@@
Several studies have explored the incorporation of cellular interaction effects into time-series scRNA-seq trajectory inference. For instance, (Atanackovic et al., 2025) employs a graph convolutional network within a flow-matching framework to model the impacts of neighborhood cells within the initial cell population. (You et al., 2024) introduced a population-level regularization in the energy form. (Yang, 2025) formulated the topological Schrödinger Bridge problem on a discrete graph. However, an explicit quantification of cell interaction dynamics in scRNA-seq trajectory inference is yet to be explored.
@@
Several works have explored mean-field problem (Zhou et al., 2024b; Ruthotto et al., 2020; Lu et al., 2024; Yang et al., 2024; Huang et al., 2024; Han et al., 2024; Shen & Wang, 2023; Li & Liu, 2025; Li et al., 2023; Shen et al., 2022) and its variants, as well as the incorporation of particle interaction terms in the Schrödinger Bridge Problem. (Backhoff et al., 2020; Hernández & Tangpi, 2025) investigated theoretical properties of the Mean Field Schrödinger Bridge Problem. (Rapakoulias et al., 2025) develop a deep learning solver for the MFSB problem. (Liu et al., 2022, 2024) proposed the generalized formulation of Schrödinger bridges that includes interacting terms. (Yang et al., 2022) formulated the ensemble regression problem and developed a neural ODE-based approach to learn the dynamics of interacting particle systems from distribution data. However, these approaches either require prior knowledge to specify the interaction potential field or do not account for unbalanced stochastic dynamics.
@@ 3.1 Regularized Unbalanced Optimal Transport
The regularized unbalanced optimal transport, also known as the unbalanced Schrödinger Bridge problem (Chen et al., 2022b), considers both the unbalanced stochastic effects in the dynamical OT framework (Baradat & Lavenant, 2021; Zhang et al., 2025a):
@@
$\inf_{\left(\rho,\mathbf{b},g\right)}\int_{0}^{1}\int_{\mathbb{R}^{d}}\frac{1}{2}\left\|\mathbf{b}(\mathbf{x},t)\right\|_{2}^{2}{\rho}(\mathbf{x},t)\mathrm{d}\mathbf{x}\mathrm{d}t+\int_{0}^{1}\int_{\mathbb{R}^{d}}\alpha\Psi\left(g(\mathbf{x},t)\right){\rho}(\mathbf{x},t)\mathrm{d}\mathbf{x}\mathrm{d}t,$
where $\Psi:\mathbb{R}\rightarrow[0,+\infty]$ corresponds to the growth penalty function, the infimum is taken over all pairs $\left(\rho,\mathbf{b},g\right)$ such that $\rho(\cdot,0)=$ $\nu_{0},\rho(\cdot,1)=\nu_{1},\rho(\mathbf{x},t)$ absolutely continuous, and
@@ 3.2 Mean Field Schrödinger Bridge Problem
$\min_{\mu^{\mathbf{X}}_{0}=\nu_{0},\leavevmode\nobreak\ \mu^{\mathbf{X}}_{1}=\nu_{1}}\mathcal{D}_{\mathrm{KL}}\left(\mu^{\mathbf{X}}_{[0,1]}\|\mu^{\mathbf{Y}}_{[0,1]}\right),$
where $\mu^{\mathbf{X}}_{[0,1]}$ is the probability measure induced by ${\mathbf{X}}_{t}\ \left(0\leq t\leq 1\right)$ and the reference measure $\mu^{\mathbf{Y}}_{[0,1]}$. However, the classical Schrödinger bridge problem considers the independent particles. The mean field extends the SB problem to the interacting particles with given initial and final distributions. We consider the $\mathrm{d}\mathbf{Y}_{t}=\bm{\sigma}(\mathbf{Y}_{t},t)\mathrm{d}{\mathbf{W}}_{t}$, where ${\mathbf{W}}_{t}\in\mathbb{R}^{d}$ and it is called the diffusion Schrödinger bridge problem, where it has a dynamic formulation. So the mean field Schrödinger bridge problem can be stated through this dynamical formulation (Backhoff et al., 2020; Hernández & Tangpi, 2025):
@@
where ϕ():+:italic-$\phi(\cdot):\mathbb{R}\rightarrow\mathbb{R}^{+}$ is the growth cost function. The infinium is taken over all $(\mathbf{b},{\rho})$ subject to $\rho(\mathbf{x},0)=\nu_{0},\rho(\mathbf{x},1)=\nu_{1}$, and
@@ 4 Unbalanced Mean Field Schrödinger Bridge
In this section, we introduce the unbalanced mean field Schrödinger Bridge problem. Inspired by regularized unbalanced optimal transport, the dynamical formulation Definition 3.2 suggests a natural way to relax the mass constraint by introducing a growth/death term $g\rho$ in (2). Meanwhile, we also introduce a loss function in (1) which considers both the growth and transition metric.
@@
In Definition 4.1, if $k(\mathbf{x},\mathbf{y})=0$, which means there is no cell-cell interaction, it degenerates to the regularized unbalanced optimal transport problem. If $\Psi({g}(\mathbf{x},t))=+\infty$ unless ${g}(\mathbf{x},t)=0$, $\Psi(0)=0$, which implies $g(\mathbf{x},t)=0$ and $k(\mathbf{x},\mathbf{y})\neq 0$, it reduces to the Mean Field Schrödinger Bridge problem, and if $k(\mathbf{x},\mathbf{y})=0$ it is the regularized optimal transport problem. It becomes the unbalanced dynamics optimal transport when $k(\mathbf{x},\mathbf{y})=\sigma(t)=0$ and $\Psi(g)=g^{2}$. It becomes the dynamics optimal transport when growth, interaction and diffusion all goes to zero. We can reformulate Definition 4.1 with the following Fisher information regularization.
@@
The unbalanced mean field Schrödinger Bridge Definition 4.1 is equivalent to
@@
where $\Psi(\cdot):\mathbb{R}\rightarrow\mathbb{R}$ is the growth cost function. The infinium is taken over all $(\mathbf{b},g,{\rho},\Phi)$ subject to $\rho(\mathbf{x},0)=\nu_{0},\rho(\mathbf{x},1)=\nu_{1}$, and
@@
Here $\mathbf{v}(\mathbf{x},t)$ is a new vector field function. This Fisher information form transforms the original SDE problem into the ODE problem which is computationally more trackable. Next, we will focus on solving this problem. The proof is simple and we left it in Section D.1 for reference.
@@
From the proof of Theorem 4.1, the relation between the new vector filed $\mathbf{v}(\mathbf{x},t)$ and $\mathbf{b}(\mathbf{x},t)$ is $\mathbf{v}(\mathbf{x},t)=\mathbf{b}(\mathbf{x},t)-\frac{1}{2}\sigma^{2}(t)\nabla_{\mathbf{x}}\log{\rho}(\mathbf{x},t).$ The new $\mathbf{v}(\mathbf{x},t)$ is also known as probability flow ODE and $\nabla_{\mathbf{x}}\log{\rho}(\mathbf{x},t)$ is the score function. Conversely, if the probability flow ODE $\mathbf{v}(\mathbf{x},t)$ and the score function $\nabla_{\mathbf{x}}\log{\rho}(\mathbf{x},t)$ are known, then we can recover the original drift term.
@@ 5 Learning Cell Dynamics and Interactions through Neural Networks
Assume that we collect scRNA-seq samples from unnormalized distributions $\mathbf{X}_{t}\in\mathbb{R}^{n_{i}\times d}$ ($t=1,2,\cdots,T$) at $T$ time points, where $n_{i}$ is the number of cells at time $i$ and $d$ is the number of genes, here we propose the CytoBridge algorithm to approximate UMFSB through neural networks. We parameterize transition velocity $\mathbf{v}(\mathbf{x},t)$, cell growth rate ${g}(\mathbf{x},t)$, log density function $\frac{1}{2}\sigma^{2}(t)\log{\rho}(\mathbf{x},t)$ and cellular interaction potential ${\Phi}(\mathbf{x},t)$ using neural networks $\mathbf{v}_{\theta},\leavevmode\nobreak\ g_{\theta}$, $s_{\theta}$ and ${\Phi}_{\theta}$ respectively. To effectively approximate the loss function in Theorem 4.1, we model the evolution of the mass densities through a number of weighted interacting particles, which is supported by the following proposition.
@@
The derivation is left in Section D.2. By combining the results in Proposition 5.1 and Theorem 4.1, the ODE constraints we simulate is indeed $\mathrm{d}\mathbf{X}_{t}^{i}=\mathbf{v}(\mathbf{X}_{t}^{i},t)\mathrm{d}t-\frac{1}{N-1}\sum_{j\neq i,j=1}^{N}{k_{i,j}w_{j}}\nabla_{\mathbf{x}}\Phi(\mathbf{X}_{t}^{i}-\mathbf{X}_{t}^{j})\mathrm{d}t$, where $k_{i,j}=k(\mathbf{X}_{t}^{i},\mathbf{X}_{t}^{j})$ is the cell-cell interaction strength and $w_{j}=w(\mathbf{X}_{t}^{j})$ is the particle weight.

### v3

@@ Abstract
Modeling the dynamics from sparsely time-resolved snapshot data is crucial for understanding complex cellular processes and behavior. Existing methods leverage optimal transport, Schrödinger bridge theory, or their variants to simultaneously infer stochastic, unbalanced dynamics from snapshot data. However, these approaches remain limited in their ability to account for cell-cell interactions. This integration is essential in real-world scenarios since intercellular communications are fundamental life processes and can influence cell state-transition dynamics. To address this challenge, we formulate the Unbalanced Mean-Field Schrödinger Bridge (UMFSB) framework to model unbalanced stochastic interaction dynamics from snapshot data. Inspired by this framework, we further propose CytoBridge, a deep learning algorithm designed to approximate the UMFSB problem. By explicitly modeling cellular transitions, proliferation, and interactions through neural networks, CytoBridge offers the flexibility to learn these processes directly from data. The effectiveness of our method has been extensively validated using both synthetic gene regulatory data and real scRNA-seq datasets. Compared to existing methods, CytoBridge identifies growth, transition, and interaction patterns, eliminates false transitions, and reconstructs the developmental landscape with greater accuracy. Code is available at: https://github.com/zhenyiizhang/CytoBridge-NeurIPS.
@@ 1 Introduction
Reconstructing dynamics from high-dimensional distribution samples is a central challenge in science and machine learning. In generative models, methods such as Variational Autoencoders (VAEs) (Kingma and Welling, 2013), diffusion models (Ho et al., 2020; Sohl-Dickstein et al., 2015; Song et al., 2021), and flow matching (Lipman et al., 2023; Tong et al., 2024a) have achieved success in generating high-fidelity images by coupling high-dimensional distributions (Liu et al., 2023). Meanwhile, in biology, inferring dynamics (also known as trajectory inference problem) from several static snapshots of single-cell RNA sequencing (scRNA-seq) data (Ding et al., 2022) to build the continuous dynamics of an individual cell and construct the corresponding cell-fate landscapes has also attracted broad interests (Schiebinger et al., 2019; Klein et al., 2025; Zhang et al., 2025b).
To study the trajectory inference problem, optimal transport (OT) theory serves as a foundational tool (Bunne et al., 2024; Zhang et al., 2025b; Heitz et al., 2024; Zhang et al., 2025c). In particular, several works propose to infer continuous cellular dynamics over time by employing the Benamou–Brenier formulation (Benamou and Brenier, 2000). Given that cell growth and death are critical biological processes, modeling the coupling of underlying unnormalized distributions has led to the development of unbalanced dynamical OT by introducing Wasserstein–Fisher–Rao metric (Chizat et al., 2018a, b). To further account for the prevalent stochastic effects on single-cell level, methods inspired by the Schrödinger Bridge (SB) problem seek to identify the most likely stochastic transition path between two arbitrary distributions (Gentil et al., 2017; Léonard, 2014). To tackle both stochastic and unbalanced effects simultaneously, methods have been developed to model unbalanced stochastic dynamics (Pariset et al., 2023; Lavenant et al., 2024), along with a recent deep learning method (Zhang et al., 2025a), which leverages the regularized unbalanced optimal transport (RUOT) framework to infer continuous unbalanced stochastic dynamics from samples without requiring prior knowledge.
@@ 2 Related Works
Numerous efforts have been made to learn dynamics from snapshot data. To tackle the dynamical optimal transport problem, several methods have been proposed by leveraging neural ODEs or flow matching techniques (Tong et al., 2020; Huguet et al., 2022; Wan et al., 2023; Zhang et al., 2024a; Tong et al., 2024a; Albergo et al., 2023; Palma et al., 2025; Rohbeck et al., 2025; Petrović et al., 2025). To account for sink and source terms in unnormalized distributions, (Peng et al., 2024; Sha et al., 2024; Tong et al., 2023; Eyring et al., 2024) developed the neural ODE-based solver for unbalanced dynamical OT. (Wang et al., 2025) developed a flow matching approach to simultaneously learn velocity and growth. For the Schrödinger Bridge (SB) problem, approaches have been proposed based on its static or dynamic formulations (Shi et al., 2024; De Bortoli et al., 2021; Gu et al., 2025; Koshizuka and Sato, 2023; Neklyudov et al., 2023, 2024; Zhang et al., 2024b; Bunne et al., 2023; Chen et al., 2022b; Zhou et al., 2024a; Zhu et al., 2024; Yeo et al., 2021; Jiang and Wan, 2024), with corresponding flow matching methods (Tong et al., 2024b). To further incorporate unbalanced effects in the SB framework, methods utilizing branching SDE theory (Lavenant et al., 2024; Ventre et al., 2023; Chizat et al., 2022), forward-backward SDE (Pariset et al., 2023), neural ODEs with Fisher information regularization (Zhang et al., 2025a), or with first-order optimality conditions (Sun et al., 2025) have been introduced. However, theoretical formulation, along with an effective deep learning solver to simultaneously account for unbalanced stochastic effects and particle interactions in the dynamical OT framework, remains largely lacking.
@@
Several studies have explored the incorporation of cellular interaction effects into time-series scRNA-seq trajectory inference. For instance, (Atanackovic et al., 2025) employs a graph convolutional network within a flow-matching framework to model the impacts of neighborhood cells within the initial cell population. (You et al., 2024) introduced a population-level regularization in the energy form. (Yang, 2025) formulated the topological Schrödinger Bridge problem on a discrete graph. (Fu et al., 2025) improves the accuracy of pseudotime inference by integrating cellular communication pattern. However, an explicit quantification of cell interaction dynamics in scRNA-seq trajectory inference is yet to be explored.
@@
Several works have explored mean-field problem (Zhou et al., 2024b; Ruthotto et al., 2020; Lu et al., 2024; Yang et al., 2024; Huang et al., 2024; Han et al., 2024; Shen and Wang, 2023; Li and Liu, 2025; Li et al., 2023; Shen et al., 2022) and its variants, as well as the incorporation of particle interaction terms in the Schrödinger Bridge Problem. (Backhoff et al., 2020; Hernández and Tangpi, 2025) investigated theoretical properties of the Mean Field Schrödinger Bridge Problem. (Rapakoulias et al., 2025) develop a deep learning solver for the MFSB problem. (Liu et al., 2022b, 2024) proposed the generalized formulation of Schrödinger bridges that includes interacting terms. (Yang et al., 2022) formulated the ensemble regression problem and developed a neural ODE-based approach to learn the dynamics of interacting particle systems from distribution data. However, these approaches either require prior knowledge to specify the interaction potential field or do not account for unbalanced stochastic dynamics.
@@ 3.1 Regularized Unbalanced Optimal Transport
The regularized unbalanced optimal transport, also known as the unbalanced Schrödinger Bridge problem (Chen et al., 2022c), considers both the unbalanced stochastic effects in the dynamical OT framework (Baradat and Lavenant, 2021; Zhang et al., 2025a):
@@
$\inf_{\left(\rho,\mathbf{b},g\right)}\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{1}{2}\left\|\mathbf{b}(\mathbf{x},t)\right\|_{2}^{2}{\rho}(\mathbf{x},t)\mathrm{d}\mathbf{x}\mathrm{d}t+\int_{0}^{T}\int_{\mathbb{R}^{d}}\alpha\Psi\left(g(\mathbf{x},t)\right){\rho}(\mathbf{x},t)\mathrm{d}\mathbf{x}\mathrm{d}t,$
where $\Psi:\mathbb{R}\rightarrow[0,+\infty]$ corresponds to the growth penalty function, and $\alpha$ is the weight of the growth penalty. The infimum is taken over all pairs $\left(\rho,\mathbf{b},g\right)$ such that $\rho(\cdot,0)=$ $\nu_{0},\rho(\cdot,1)=\nu_{1},\rho(\mathbf{x},t)$ absolutely continuous, and
@@ 3.2 Mean Field Schrödinger Bridge Problem
$\min_{\mu^{\mathbf{X}}_{0}=\nu_{0},~\mu^{\mathbf{X}}_{1}=\nu_{1}}\mathcal{D}_{\mathrm{KL}}\left(\mu^{\mathbf{X}}_{[0,1]}\|\mu^{\mathbf{Y}}_{[0,1]}\right),$
where $\mu^{\mathbf{X}}_{[0,1]}$ is the probability measure induced by ${\mathbf{X}}_{t}\ \left(0\leq t\leq 1\right)$ and the reference measure $\mu^{\mathbf{Y}}_{[0,1]}$. However, the classical Schrödinger bridge problem considers the independent particles. The mean field extends the SB problem to the interacting particles with given initial and final distributions. We consider the $\mathrm{d}\mathbf{Y}_{t}=\bm{\sigma}(\mathbf{Y}_{t},t)\mathrm{d}{\mathbf{W}}_{t}$, where $\bm{\sigma}(\mathbf{Y}_{t},t)\in\mathbb{R}^{d\times d}$ is the diffusion rate, ${\mathbf{W}}_{t}\in\mathbb{R}^{d}$ is the standard multi-dimensional Brownian motion and it is called the diffusion Schrödinger bridge problem, where it has a dynamic formulation. So the mean field Schrödinger bridge problem can be stated through this dynamical formulation (Backhoff et al., 2020; Hernández and Tangpi, 2025):
@@
The infinium is taken over all $(\mathbf{b},{\rho})$ subject to $\rho(\mathbf{x},0)=\nu_{0},\rho(\mathbf{x},1)=\nu_{1}$, and
@@ 4 Unbalanced Mean Field Schrödinger Bridge
In this section, we introduce the unbalanced mean field Schrödinger Bridge problem. Inspired by regularized unbalanced optimal transport, the dynamical formulation Definition˜3.2 suggests a natural way to relax the mass constraint by introducing a growth/death term $g\rho$ in (2). Meanwhile, we also introduce a loss function in (1) which considers both the growth and transition metric.
@@
In Definition˜4.1, if $k(\mathbf{x},\mathbf{y})=0$, which means there is no cell-cell interaction, it degenerates to the regularized unbalanced optimal transport problem. If the growth penalty is set such that ${g}(\mathbf{x},t)$ must be zero (i.e., by setting $\Psi({g}(\mathbf{x},t))=+\infty$ for ${g}(\mathbf{x},t)=0$ and $\Psi(0)=0$), the framework degenerates to simpler forms: if interactions are present ($k(\mathbf{x},\mathbf{y})\neq 0$), the formulation reduces to the Mean Field Schrödinger Bridge problem; if interactions are absent ($k(\mathbf{x},\mathbf{y})=0$), it reduces to the regularized optimal transport problem. It becomes the unbalanced dynamics optimal transport when $k(\mathbf{x},\mathbf{y})=\sigma(t)=0$ and $\Psi(g)=g^{2}$. It becomes the dynamics optimal transport when growth, interaction and diffusion all goes to zero. We can reformulate Definition˜4.1 with the following Fisher information regularization.
@@
The unbalanced mean field Schrödinger Bridge Definition˜4.1 is equivalent to
@@
where $\Psi(\cdot):\mathbb{R}\rightarrow\mathbb{R}$ is the growth cost function, and $\alpha$ is the weight of the growth cost. The infinium is taken over all $(\mathbf{b},g,{\rho},\Phi)$ subject to $\rho(\mathbf{x},0)=\nu_{0},\rho(\mathbf{x},1)=\nu_{1}$, and
@@
Here $\mathbf{v}(\mathbf{x},t)$ is a new vector field function. This Fisher information form transforms the original SDE problem into the ODE problem which is computationally more tractable. Next, we will focus on solving this problem. The proof is simple and we left it in Section˜D.1 for reference.
@@
From the proof of Theorem˜4.1, the relation between the new vector filed $\mathbf{v}(\mathbf{x},t)$ and $\mathbf{b}(\mathbf{x},t)$ is $\mathbf{v}(\mathbf{x},t)=\mathbf{b}(\mathbf{x},t)-\frac{1}{2}\sigma^{2}(t)\nabla_{\mathbf{x}}\log{\rho}(\mathbf{x},t).$ The new $\mathbf{v}(\mathbf{x},t)$ is also known as probability flow ODE and $\nabla_{\mathbf{x}}\log{\rho}(\mathbf{x},t)$ is the score function. Conversely, if the probability flow ODE $\mathbf{v}(\mathbf{x},t)$ and the score function $\nabla_{\mathbf{x}}\log{\rho}(\mathbf{x},t)$ are known, then we can recover the original drift term.
@@ 5 Learning Cell Dynamics and Interactions through Neural Networks
Assume that we collect scRNA-seq samples from unnormalized distributions $\mathbf{X}_{t}\in\mathbb{R}^{n_{i}\times d}$ ($t=1,2,\cdots,T$) at $T$ time points, where $n_{i}$ is the number of cells at time $i$ and $d$ is the number of genes, here we propose the CytoBridge algorithm to approximate UMFSB through neural networks. We parameterize transition velocity $\mathbf{v}(\mathbf{x},t)$, cell growth rate ${g}(\mathbf{x},t)$, log density function $\frac{1}{2}\sigma^{2}(t)\log{\rho}(\mathbf{x},t)$ and cellular interaction potential ${\Phi}(\mathbf{x},t)$ using neural networks $\mathbf{v}_{\theta},~g_{\theta}$, $s_{\theta}$ and ${\Phi}_{\theta}$ respectively, as shown in Fig.˜1. To effectively approximate the loss function in Theorem˜4.1, we model the evolution of the mass densities through a number of weighted interacting particles, which is supported by the following proposition.
@@
The derivation is left in Section˜D.2. By combining the results in Proposition˜5.1 and Theorem˜4.1, the ODE constraints we simulate is indeed $\mathrm{d}\mathbf{X}_{t}^{i}=\mathbf{v}(\mathbf{X}_{t}^{i},t)\mathrm{d}t-\frac{1}{N-1}\sum_{j\neq i,j=1}^{N}{k_{i,j}w_{j}}\nabla_{\mathbf{x}}\Phi(\mathbf{X}_{t}^{i}-\mathbf{X}_{t}^{j})\mathrm{d}t$, where $k_{i,j}=k(\mathbf{X}_{t}^{i},\mathbf{X}_{t}^{j})$ is the cell-cell interaction strength and $w_{j}=w(\mathbf{X}_{t}^{j})$ is the particle weight.